Using the FACTORIAL RULE, respond to the following questions. Each of these distinct questions is asking you to count the total number or combinations using this rule. How many ways can 7 paintings be lined up [Choose ] on a wall? How many 5-number license plates can be Choose ] made using the digits 0, 2, 4, 6, 8 if repetitions are NOT allowed? How many 4-digit numbers can be formed ifChoose each one uses all the digits 0, 1, 2, 3, 4, 5, 6, 7 where repetitions are NOT allowed? 10 different books are on a shelf. In how Choose ] many different ways can you arrange them?
Answer #1

Answer - 1: Number of ways 7 paintings can be lined up = 7! = 7*6*5*4*3*2*1 = 5040

Answer - 2: Number of license plates = 4*4*3*2*1 = 96

Answer - 3: Number of 4 - digit numbers = 6*6*5*4 = 720

Answer - 4: Number of ways = 10! = 10*9*8*7*6*5*4*3*2*1 = 3628800